OpenStudy (emac):

Find the Derivative!! -7x^4sqrtx+-7/x^3sqrtx

OpenStudy (518nad):

sqrt x = x^0.5

OpenStudy (518nad):

can you simplify this x^a*x^b = x^(a+b) you can use this

OpenStudy (emac):

When I changed sqrtx to x^1/2 and combined with the other x's I got -7x^4.5-7x^-3.5

OpenStudy (legomyego180):

\[-7x^4\sqrt{x}+\frac{ -7 }{ x^3\sqrt{x} }\] Is this right?